Calculate the standard enthropy change to produce synthesis gas




CO(g) + H2(g) -> CH4(g) + H2O(g)

1
Expert's answer
2021-10-30T01:13:24-0400

CO(g) + 3H2(g) → CH4(g) + H2O(g)

We know the ΔH0:


"ΔH^0 = \sum ΔH^0_{products} -\sum ΔH^0_{reactants} \\

= [ΔH^0 CH_4 (g) + ΔH^0 H_2O (g)] – [ΔH^0 H_2 (g) + ΔH^0 CO(g)] \\

= [-74.6 +(-241.8)] -[3 \times 0 + (-110.5)] \\

= -316.4 +110.5 \\

= -205.9 \;kJ"